# Measurement of the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v1

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We show that these observables are more sensitive to the transport coefficients than the traditional flow observables.
@@ I Introduction
In ultra-relativistic heavy-ion collisions, a thermalized, strongly interacting, deconfined medium, known as Quark-Gluon Plasma (QGP), is expected to form [1, 2, 3, 4, 5, 6]. The information about such an event is carried both by anisotropies in the initial state collision geometry and by the transport properties of QGP [7, 8, 3]. After hadronization, particles are emitted anisotropically in the transverse plane to the beam direction, as detected by a detector. Traditionally, Fourier series in flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ are used to quantify this azimuthal distribution in the momentum space [9],
@@
Collective flow is crucial for studying the properties of the QGP medium formed in these collisions [10, 7, 11]. The impact parameter driven spatial geometry of the fireball, event-by-event nucleon position as well as sub-nucleonic fluctuations within the overlap region of the two colliding nuclei lead to the development of various orders of anisotropic flow harmonics. Due to the almond shape of the overlap region, the primary source of the second-order flow harmonic, known as elliptic flow ($v_{2}$), is the initial spatial anisotropy [12, 13, 14, 15, 16, 17, 18]. The third-order flow harmonic, known as triangular flow ($v_{3}$), arises from event-by-event fluctuations in the positions of the colliding nucleons as well as their sub-nucleonic fluctuations [19, 20, 21, 22, 23]. These fluctuations also give rise to higher-order flow harmonics like $v_{4}$, $v_{5}$, etc. Previous studies have demonstrated that these flow harmonics are sensitive to the equation of state (EoS) and the transport properties of the fireball created in the collision, such as shear and bulk viscosity [24, 25, 26, 27, 28].
Anisotropic flow analysis involves measuring $v_{n}$, $\psi_{n}$ and their event-by-event correlations and fluctuations. Due to the random fluctuations of the impact parameter vector, however, the estimation of $v_{n}$ and $\psi_{n}$ involves the complications associated with reaction plane dispersion in conventional flow analyses. As a result, they are indirectly estimated using correlation techniques, in which there is no need for an event-by-event determination of the reaction plane [29, 30]. Therefore, there is no need to correct for dispersion in an estimated reaction plane. The foundation of this alternative approach is based on the following outcome,
@@
The double angular brackets indicate that the averaging procedure is performed in two steps—first, averaging over all distinct particle multiplets in an event and then averaging these single-event averages with appropriately chosen event weights in the second step. By generalizing this concept for non-identical harmonics, new observables, which strictly satisfy all defining mathematical properties of cumulants, were constructed to quantify the correlations among different flow harmonics [34]. Ref. [35] further generalized these observables to probe the genuine correlation between different moments of flow harmonics.
@@ I.1 Symmetric Cumulants of Flow Amplitudes
Reference [34] proposed a general algorithm to measure multi-particle correlation where the harmonics inside the correlator bracket of Eq. (3) can be different. This introduces a new set of observables known as symmetric cumulants $(SCs)$ which can be used to measure the correlation between event-by-event fluctuation of flow harmonics $v_{m}$ and $v_{n}$. These four-particle symmetric cumulants, $SC(m,n)$, are defined as [36]
$SC(m,n)$$\equiv$ $\cos(m\varphi_{1}\!+\!n\varphi_{2}\!-\!m\varphi_{3}-\!n\varphi_{4})\right>\right>_{c}$ (4)
@@
The subscript $c$ indicates the cumulant. Positive (Negative) values of $SC(m,n)$ suggest the correlation (anti-correlation) between $v_{m}^{2}$ and $v_{n}^{2}$, which means that if $v_{m}^{2}$ is larger than $\langle v_{m}^{2}\rangle$ in an event then the probability of $v_{n}^{2}$ being larger than $\langle v_{n}^{2}\rangle$ in that same event is enhanced (suppressed). The $SC(m,n)$ observables focus on the correlations among different orders of flow harmonics and enable the quantitative comparison between experimental data and model calculations. To eliminate the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant, we divide $SC(m,n)$ by their average values, $v_{m}^{2}\right>$ and $v_{n}^{2}\right>$, and define the normalized symmetric cumulant. This enables us to compare data and model calculations in a quantitative way, and to compare the fluctuations of the initial and final states. The normalized symmetric cumulant, denoted by $NSC(m,n)$, is achieved following the standard method from Ref. [37]:
$NSC(m,n)=\frac{SC(m,n)}{\langle v_{m}^{2}\rangle\langle v_{n}^{2}\rangle}.$ (5)
These correlations have been measured by both STAR experiment at RHIC [38] and ALICE experiment at LHC [39] with observation of a positive correlation between $v_{2}$ and $v_{4}$ while a negative correlation between $v_{2}$ and $v_{3}$. The sensitivity of these observables to the shear viscosity to entropy density ($\eta/s$) has been observed in both hydrodynamics and transport model studies  [40, 41].
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
Recent studies indicate that insightful information can be obtained about the properties of Quark-Gluon Plasma by using higher-order observables [42]. These observables can probe the genuine correlations between the different moments of different flow harmonics. They are robust against non-flow correlations which can be verified using the HIJING Monte Carlo generator [43]. The asymmetric cumulant ${\rm AC}_{2,1}(m,n)$ is defined as [35]
@@
In Eq. (6), the subscript (2,1) on the left-hand side indicates the exponents of $v_{m}^{2}$ and $v_{n}^{2}$ on the right-hand side. The final combinations of azimuthal correlators used to estimate the $ACs$ experimentally are [35]:
@@
These expressions are genuine multivariate cumulants. ${\rm AC}_{1,1}(m,n)$ corresponds to SC$(m,n)$, illustrating the generalisation aspect of the $ACs$. We can also normalise the $ACs$. This procedure has two benefits. Normalizing the results allows proper comparisons and determination of the initial state effects and the changes brought by the hydrodynamic evolution since the predictions for $ACs$ in the initial and final state do not have the same scale. Flow amplitudes depend on the transverse momentum, $p_{T}$, which leads to a similar dependence in any linear combinations of them, such as the $SCs$ and the $ACs$. The normalization eliminates this dependence and enables comparisons between models and data with different $p_{T}$ ranges [43]. Again, the normalization of the $ACs$ is done following the standard method from Ref. [37],
@@
In this paper, we have presented the measurement of $SC(m,n)$ and $AC_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. $SC(m,n)$ have been measured in previous studies from hydrodynamic and transport models [44, 43, 37, 45, 46, 47, 48], while $AC_{2,1}(m,n)$ have been measured in transport models [46, 47]. ALICE collaboration has measured $AC_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [49]. In addition to symmetric cumulants, various other correlators have also been studied [50, 51, 52, 53]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
A hyper-surface is generated from the hydrodynamic space-time evolution of the fluid. To describe the dilute hadronic phase, we utilize the iSS code [65, 66] to sample the primordial hadrons from the hypersurface with a constant energy density, $e_{sw}=0.26\ GeV/fm^{3}$, equivalent to a local temperature of approximately 151 MeV. Then we use the UrQMD code [67, 68] to simulate scatterings and decays of hadrons during the late stage of heavy ion collisions. To increase statistics, each hydrodynamic switching hyper-surface is sampled multiple times. The number of oversampling events for every hydrodynamic event is determined to ensure sufficient statistics for every centrality class. We generated ensembles of 2000 hydrodynamic events for each centrality class, each with a suitable number of repeated samplings. Table 1 lists the number of events generated for each of the centrality classes with the three sets of parameters for the hydrodynamic stage shown in Table 2:
@@
Table 1: Ensemble size (in millions) of the centrality classes.
@@
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 3 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model.
@@ III Results
We calibrated our model by using the produced particle yields, transverse momentum spectra and elliptic flow. We then compared these results to experimental data from the PHENIX and STAR Collaborations. The normalization factor for the system’s energy density was determined by matching the model and STAR data [69] charged hadron multiplicity $dN_{Ch}/d\eta$ in the 0-5% centrality bin. We also compared the $p_{T}$ spectra of identified particles in 0-5% centrality Au+Au collisions to STAR results [69]. Additionally, we compared the single-particle anisotropic flow coefficient $v_{2}$ of charged particles in 20-30% centrality Au+Au collisions with the PHENIX measurement [70]. The hydrodynamic evolution, of the MCGM initial conditions, coupled to UrQMD can reproduce the elliptic flow upto the mid-central collisions.
@@ III.1.1 Four-particle symmetric cumulants
In this section, we present results for $SC(m,n)$ from our simulations. To check their sensitivity to the hydro transport coefficients, we compute these observables for the three sets of ensembles mentioned in Table  2. We observe that both $SC(2,3)$ and $SC(2,4)$ are suppressed by shear and bulk viscosities. In Fig.[1], we show the effect of shear viscosity on $v_{2}\{4\}$ and $SC(2,3)$. In $v_{2}\{4\}$, we see a maximum change of around 10% in the mid-central collisions. While in $SC(2,3)$, we see a change of around 40-60% throughout the centrality range. We observe a similar dependence on the bulk viscosity of the medium. Thus, these observables are more sensitive to the transport coefficients. Similar to $SC(2,3)$, $SC(2,4)$ is also affected by the transport coefficients in the same manner. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals a similar difference in their sensitivity to shear and bulk viscosity.
Figure 1: Ratios of $v_{2}\{4\}$ and $SC(2,3)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and $SC(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and $SC(2,3)$ using markers. Certain data points with high uncertainty were excluded for clarity.
@@ III.1.2 Six-particle asymmetric cumulants
Now, we present the results for $AC_{2,1}(m,n)$. In Fig.[2], we show the effect of shear viscosity on elliptic flow from the four-particle correlations and $AC_{2,1}(2,3)$. For $AC_{2,1}(2,3)$, we see a change of around 50% throughout the centrality range. We observe that the medium’s bulk viscosity also has a similar effect on these observables. Compared to the traditional flow observables, we see that these are more sensitive to the transport coefficients. Hence, these will put better constraints on the transport properties of the medium created in heavy-ion collisions.
Figure 2: Ratios of $v_{2}\{4\}$ and $AC_{2,1}(2,3)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and $AC_{2,1}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and $AC_{2,1}(2,3)$ using markers. Certain data points with high uncertainty were excluded for clarity.
III.2 Centrality dependence of $SC(m,n)$ and $AC_{2,1}(m,n)$

### v2

# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We show that these observables are more sensitive to the transport properties than the traditional flow observables.
@@ I Introduction
In ultra-relativistic heavy-ion collisions, a thermalized, strongly interacting, deconfined medium, known as Quark-Gluon Plasma (QGP), is expected to form [1, 2, 3, 4, 5, 6]. The information regarding certain properties of QGP is carried by anisotropies observed in the momentum distributions of the produced particles [7, 8, 3]. After hadronization, particles are emitted anisotropically in the transverse plane to the beam direction, as detected by a detector. Traditionally, Fourier series in flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ are used to quantify this azimuthal distribution in the momentum space [9],
@@
Collective flow is crucial for studying the properties of the QGP medium formed in these collisions [10, 7, 11]. The impact parameter driven spatial geometry of the fireball, event-by-event nucleon position as well as sub-nucleonic fluctuations within the overlap region of the two colliding nuclei lead to the development of various orders of anisotropic flow harmonics. Due to the almond shape of the overlap region, the primary source of the second-order flow harmonic, known as elliptic flow ($v_{2}$), is the initial spatial anisotropy [12, 13, 14, 15, 16, 17, 18]. The third-order flow harmonic, known as triangular flow ($v_{3}$), arises from event-by-event fluctuations in the positions of the colliding nucleons as well as their sub-nucleonic fluctuations [19, 20, 21, 22, 23]. These fluctuations also give rise to higher-order flow harmonics. Previous studies have demonstrated that these flow harmonics are sensitive to the equation of state (EoS) and the transport properties of the fireball created in the collision, such as shear and bulk viscosity [24, 25, 26, 27, 28].
Anisotropic flow analysis involves measuring $v_{n}$, $\psi_{n}$, and their event-by-event correlations and fluctuations. Due to the random fluctuations of the impact parameter vector, however, the estimation of $v_{n}$ and $\psi_{n}$ involves the complications associated with reaction plane dispersion in conventional flow analyses. As a result, they are indirectly estimated using correlation techniques, in which there is no need for an event-by-event determination of the reaction plane [29, 30]. Therefore, there is no need to correct for dispersion in an estimated reaction plane. The foundation of this alternative approach is based on the following outcome,
@@
The double angular brackets indicate that the averaging procedure is performed in two steps, averaging over all distinct particle multiplets in an event and then averaging these single-event averages with appropriately chosen event weights. By generalizing this concept for non-identical harmonics, new observables, which strictly satisfy all defining mathematical properties of cumulants, were constructed to quantify the correlations among different flow harmonics [34]. Ref. [35] further generalized these observables to probe the genuine correlation between different moments of flow harmonics.
@@ I.1 Symmetric Cumulants of Flow Amplitudes
Reference [34] proposed a general algorithm to measure multi-particle correlation where the harmonics inside the correlator bracket of Eq. (3) can be different. This introduces a new set of observables, known as symmetric cumulants $(SCs)$, which can be used to measure the correlation between event-by-event fluctuation of flow harmonics $v_{m}$ and $v_{n}$. This new approach allows for the separation of non-flow and flow contributions and provides initial insights into how the combinatorial background contributes to flow measurements using correlation techniques. These four-particle symmetric cumulants, ${\rm SC}(m,n)$, are defined as [36]
${\rm SC}(m,n)$$\equiv$ $\cos(m\varphi_{1}\!+\!n\varphi_{2}\!-\!m\varphi_{3}-\!n\varphi_{4})\right>\right>_{c}$ (4)
@@
The subscript $c$ indicates the cumulant. Positive (Negative) values of ${\rm SC}(m,n)$ suggest the correlation (anti-correlation) between $v_{m}^{2}$ and $v_{n}^{2}$, which means that if $v_{m}^{2}$ is larger than $\langle v_{m}^{2}\rangle$ in an event then the probability of $v_{n}^{2}$ being larger than $\langle v_{n}^{2}\rangle$ in that same event is enhanced (suppressed). The ${\rm SC}(m,n)$ observables focus on the correlations among different orders of flow harmonics and enable the quantitative comparison between experimental data and model calculations. To eliminate the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant, we divide ${\rm SC}(m,n)$ by their average values, $v_{m}^{2}\right>$ and $v_{n}^{2}\right>$, and define the normalized symmetric cumulant. This enables us to compare data and model calculations in a quantitative way and compare the fluctuations of the initial and final states. The normalized symmetric cumulant, denoted by ${\rm NSC}(m,n)$, is achieved following the standard method from Ref. [37]:
${\rm NSC}(m,n)=\frac{{\rm SC}(m,n)}{\langle v_{m}^{2}\rangle\langle v_{n}^{2}\rangle}.$ (5)
These correlations have been measured by both STAR experiment at RHIC [38] and ALICE experiment at LHC [39] with the observation of a positive correlation between $v_{2}$ and $v_{4}$ while a negative correlation between $v_{2}$ and $v_{3}$. The sensitivity of these observables to the shear viscosity to entropy density ($\eta/s$) has been studied in transport models like AMPT [40].
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
Recent studies indicate that insightful information can be obtained about the properties of Quark-Gluon Plasma by using higher-order observables [41]. These observables can probe the genuine correlations between the different moments of different flow harmonics. They are robust against non-flow correlations, which can be verified using the HIJING Monte Carlo generator [42]. The asymmetric cumulant ${\rm AC}_{2,1}(m,n)$ is defined as [35]
@@
In Eq. (6), the subscript (2,1) on the left-hand side indicates the exponents of $v_{m}^{2}$ and $v_{n}^{2}$ on the right-hand side. The combinations of azimuthal correlators used to estimate the $ACs$ are [35]:
@@
These expressions are genuine multivariate cumulants. ${\rm AC}_{1,1}(m,n)$ corresponds to SC$(m,n)$, illustrating the generalisation aspect of the $ACs$. We can also normalize the $ACs$. This procedure has two benefits. Normalizing the results allows proper comparisons and determination of the initial state effects and the changes brought by the hydrodynamic evolution since the predictions for $ACs$ in the initial and final state do not have the same scale. Flow amplitudes depend on the transverse momentum, $p_{T}$, which leads to a similar dependence in any linear combinations of them, such as the $SCs$ and the $ACs$. The normalization eliminates this dependence and enables comparisons between models and data with different $p_{T}$ ranges [42]. Again, the normalization of the $ACs$ is done following the standard method from Ref. [37],
@@
In this paper, we have presented the measurement of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 42, 37, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [45, 46]. ALICE collaboration has measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In addition to symmetric cumulants, various other correlators have also been studied [49, 50, 51, 52]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
A hyper-surface is generated from the hydrodynamic space-time evolution of the fluid. To describe the dilute hadronic phase, we utilize the iSS code [64, 65] to sample the primordial hadrons from the hypersurface with a constant energy density, $e_{sw}=0.26\ GeV/fm^{3}$, equivalent to a local temperature of approximately 151 MeV. Then, we use the UrQMD code [66, 67] to simulate scatterings and decays of hadrons during the late stage of heavy ion collisions. To increase statistics, each hydrodynamic switching hyper-surface is sampled multiple times. The number of oversampling events for every hydrodynamic event is determined to ensure sufficient statistics for every centrality class. We generated ensembles of 2000 hydrodynamic events for each centrality class, each with a suitable number of repeated samplings. Table 1 lists the number of events generated for each of the centrality classes with the three sets of parameters for the hydrodynamic stage shown in Table 2. Table 3 outlines the particles used in the analyses at the end of different stages of the evolution. Analysis at the end of:
$\bullet$ stage I includes hydrodynamic evolution and particlization,
$\bullet$ stage II includes hydrodynamic evolution, particlization, and resonance decay, and
$\bullet$ stage III includes hydrodynamic evolution, particlization, resonance decay, and hadronic interactions within the UrQMD model approach.
@@
Table 1: Ensemble size (in millions) of the centrality classes at the end of Stage III.
@@
Particle list
Stage I Hydrodynamic evolution + particlization (Primordial)
Stage II Primordial + Resonace decays
Stage III Primordial + UrQMD
Table 3: The particle lists analyzed at the end of different stages of the evolution (Hydrodynamic evolution, resonance decay, and hadronic transport).
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
@@ III Results
We calibrated our model by using the produced particle yields, transverse momentum spectra and elliptic flow. We then compared these results to experimental data from the PHENIX and STAR Collaborations. The normalization factor for the system’s energy density was determined by matching the model and STAR data [68] charged hadron multiplicity $dN_{Ch}/d\eta$ in the 0-5% centrality bin. We also compared the $p_{T}$ spectra of identified particles in 0-5% centrality Au+Au collisions to STAR results [68]. Additionally, we compared the single-particle anisotropic flow coefficient $v_{2}$ of charged particles in 20-30% centrality Au+Au collisions with the PHENIX measurement [69]. The hydrodynamic evolution of the MCGM initial conditions, coupled to UrQMD, can reproduce the elliptic flow up to the mid-central collisions.
@@ III.1.1 Four-particle symmetric cumulants
In this section, we present results for ${\rm SC}(m,n)$ from our simulations. To check their sensitivity to the hydro transport coefficients, we compute these observables for the three sets of ensembles mentioned in Table  2. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by shear and bulk viscosities. In Fig.[1], we show the effect of shear viscosity on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a maximum change of around 10% in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40-50% throughout the centrality range. We observe a similar dependence on the bulk viscosity of the medium. Thus, these observables are more sensitive to the transport coefficients. Similar to ${\rm SC}(2,3)$, ${\rm SC}(2,4)$ is also affected by the transport coefficients in the same manner. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals a similar difference in their sensitivity to shear and bulk viscosity.
Figure 1: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
@@ III.1.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[2], we show the effect of shear viscosity on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 50% throughout the centrality range (note that 0-10% and 10-20% centrality classes have large statistical uncertainties). We observe that the medium’s bulk viscosity also has a similar effect on these observables. Compared to the traditional flow observables, we see that these are more sensitive to the transport coefficients.
Figure 2: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
III.2 Sensitivity to hadronic interactions